OpenStudy (jhannybean):

Evaluate double integral 4y^3 dA bound by y = x-6, y^2 = x

OpenStudy (kainui):

If you can't visualize it, then you gotta draw a picture.

OpenStudy (swissgirl):



OpenStudy (jhannybean):

the intersections are what are messing me up T_T

OpenStudy (swissgirl):

well see where they intersect its a line with a parabola

OpenStudy (jhannybean):

is it just.... y^2-y-6 = 0 (y-3)(y+2) = 0 y 3, -2?